OpenStudy (anonymous):

solve for -2(2x-5)=2

OpenStudy (anonymous):

Do you know what the first step is?

OpenStudy (anonymous):

um no i think its just solve for x

OpenStudy (anonymous):

you have to distribute the -2 to 2x and -5 making the equation -4x+10=2 do you know what to do from there?

OpenStudy (anonymous):

-4x+10=2 -4x+10-10=2-10 -4x=-8 -4x/-4=-8/-4 x=2

OpenStudy (anonymous):

oh okay i get it thank yu.